OpenStudy (anonymous):

Mark and Nancy are making punch for a party and are using two types of juice: pineapple and papaya. The total number of quarts must be more than 9 but less than 15. They should use at least 4 quarts of each type. We want to find out how many quarts of each juice they should buy. Pick one ordered pair solution for this problem and tell what each number represents. Mark and Nancy are making punch for a party and are using two types of juice: pineapple and papaya. The total number of quarts must be more than 9 but less than 15. They should use at least 4 quarts of each type. We want to find out how many quarts of each juice they should buy. Pick one ordered pair solution for this problem and tell what each number represents. @Mathematics

OpenStudy (turingtest):

x=quarts of papaya y=quarts of pineapple let x=5 and y=6 (5,6) is in the solution set because \[5+6=11\ge9\]\[5+6=11\le15\]\[5\ge4\]\[6\ge4\] so 5 quarts of papaya and 6 quarts pf pineapple satisfies all our inequalities.